OpenStudy (anonymous):

Simplify. Do not use negative exponents in the answer. y^9/y^-1 would = y^10/y?

OpenStudy (anonymous):

o, zero

OpenStudy (anonymous):

y^10

OpenStudy (anonymous):

the y from the bottom would move up making it y^10/1 or y^10

OpenStudy (anonymous):

\[y^{9}/y^{-1}=y^{9}/\frac{1}{y}=y^{9}*y=y^{10}\]

OpenStudy (anonymous):

i feel person has asked what will be at the question mark place i.e., y^9/y^(-1) = y^10 / y( ? )

OpenStudy (anonymous):

so at the place of question mark it wud be ZERO, 0